OpenStudy (anonymous):

find the coordinates of K, the midpoint of of MN, if the endpoints of MN are M(-4,6) and N (2,5) . Answer (0, 4 2/3) (-1, 5 ½) (-2, 6) (3, 5 ½)

OpenStudy (anonymous):

(-1, 5 ½)

OpenStudy (anonymous):

i.e ((-4+2)/2,(5+6)/2)

OpenStudy (lalaly):

midpoint of MN x=(2+-4 )/2 , y=5+6/2 x=-1 y=5.5 so second answer